# Experiential Semantic Information and Brain Alignment: Are Multimodal Models Better than Language Models?

### v1

@@ Abstract
A common assumption in Computational Linguistics is that text representations learnt by multimodal models are richer and more human-like than those by language-only models, as they are grounded in images or audio—similar to how human language is grounded in real-world experiences. However, empirical studies checking whether this is true are largely lacking. We address this gap by comparing word representations from contrastive multimodal models vs. language only ones in the extent to which they capture experiential information—as defined by an existing norm-based ‘experiential model’—and align with human fMRI responses. Our results indicate that, surprisingly, language-only models are superior to multimodal ones in both respects. Additionally, they learn more unique brain-relevant semantic information beyond that shared with the experiential model. Overall, our study highlights the need to develop computational models that better integrate the complementary semantic information provided by multimodal data sources.
@@ 1 Introduction
The relationship between abstract linguistic representations and the real-world entities they refer to is central to the academic discourse around semantics—the ‘symbol-grounding problem’ Harnad (1990); Bender and Koller (2020). While some researchers think that word meanings should be intended as purely symbolic Fodor (1983), others believe that words have meanings precisely because they are linked to specific entities, experiences or notions Barsalou (2008). Supporters of the latter view stress that human language acquisition is situated in a rich multimodal environment, where new words are learnt through interactions with objects and people Vigliocco et al. (2014). Theories of embodied cognition further highlight the importance of linking words to concrete experience not only for their acquisition but also for their comprehension. Indeed, according to these theories, understanding sentences involves engaging perceptual, motor or emotional simulations of their content (for an overview, see Kaschak et al., 2024).
@@
Operationalising the ‘extra-linguistic’ information that multimodal models are allegedly learning is a prerequisite for approaching this issue. We did this by relying on a semantic model introduced by Fernandino et al. (2022) to capture ‘experiential information’. This cognitive model represents words as $n$-dimensional arrays where each entry corresponds to aggregated human ratings on a pre-defined experiential attribute (e.g., Vision, Motion, Harm). We focused on a set of nouns and evaluated the alignment between their representations provided by the experiential model and those by comparable unimodal (language-only) and multimodal (vision-language and audio-language) computational models. This analysis allowed us to uncover if multimodal models indeed reflect more semantic information than language-only models. Next, we checked whether capturing experiential information translates into higher alignment with fMRI brain responses to the same set of noun words.
Our findings indicate several interesting trends. First, both vision-language and language-only models exhibit significant alignment with the experiential model and brain responses, while the audio-language model displays weak or non-significant correlations. Second, this alignment is more pronounced for language-only models, which appear to capture a great deal of brain-relevant information beyond experiential. Lastly, language-only models remain more brain-aligned than vision-language models even when focusing on a set of more concrete words, although the gap is reduced. Overall, our study shows that current multimodal models learn less brain-relevant information—both experiential and beyond—than comparable language-only models, highlighting the need to explore different approaches to construct multimodal word representations. The code to reproduce our experiments will be released upon publication.
@@ 2.1 Embodied cognition
For example, a series of works on the Action-sentence Compatibility Effect (ACE, Glenberg and Kaschak, 2002) and its subsequent variants Borreggine and Kaschak (2006); Zwaan and Taylor (2006); Bub and Masson (2012) revealed a significant difference in reaction times—attributed to motor simulations—when participants had to respond to a sentence (e.g., You passed the note to Art) with a movement (extending their arm) matching vs.  non-matching (retreating their arm) that mentioned in the sentence. Similarly, the sentence-picture verification task Stanfield and Zwaan (2001), where participants have to respond to a picture that is either compatible (an eagle with its wings outstretched) or incompatible (an eagle with its wings folded) with a sentence (The eagle is in the sky), and its variations Connell (2007); Hoeben Mannaert et al. (2017) have also been widely used to demonstrate the occurrence of perceptual simulation during language comprehension. In parallel, a line of neuroscientific studies have found evidence that semantic processing may activate motor (among others, Hauk et al., 2004; Tettamanti et al., 2005; Aziz-Zadeh et al., 2006) and perceptual brain regions Kiefer et al. (2008); Van Dam et al. (2012).
@@ 2.2 Multimodal models of semantics
Embodied cognition and related ideas, such as visual grounding, have percolated from Cognitive Science to Computational Linguistics, motivating attempts to build semantic models that learn representations from data sources beyond text. Early efforts in this direction (e.g., Bruni et al., 2014; Kiela and Bottou, 2014; Lazaridou et al., 2015; Silberer and Lapata, 2012, 2014) were characterised by 1) a focus on developing human-aligned computational models of meaning and 2) limited computational modelling resources available (large datasets of paired image-text inputs did not exist at the time, nor did large transformer-based architectures).
Recently, multimodal models have become more powerful and found application on a variety of downstream tasks (e.g., image captioning, image retrieval, or visual question answering). Some seminal works used a contrastive objective to learn aligned image and text representations Radford et al. (2021); Jia et al. (2021), while others—often inspired by BERT’s Devlin et al. (2019) successes in language modelling—tried to apply its underlying intuitions to the vision-language domain Tan and Bansal (2019); Li et al. (2019); Lu et al. (2019); Chen et al. (2020). Finally, state-of-the-art large vision-language models (Li et al., 2023; Wang et al., 2024; Liu et al., 2024; Deitke et al., 2024; Laurençon et al., 2024, LVLMs, e.g.,), which usually bridge together a large language model (LLM) and an image encoder, can engage in strikingly human-like conversations about images. In contrast to the early attempts at multimodal modelling, these works share 1) a focus on solving, or improving performance on, downstream tasks, and 2) the availability of massive datasets and large models with billions of parameters.
@@ 2.3 Experiential models of semantics
Recently, a few approaches motivated by embodied cognition have introduced models of semantics aimed at capturing ‘experiential information’, i.e., aspects of meaning related to how humans ground language in experiences. These experiential models were constructed by asking human annotators to rank words on a set of pre-defined dimensions. For example, Fernandino et al. (2022) introduced an experiential model based on 48 dimensions spanning perceptual, emotional, and action-related constructs. In two fMRI studies, they found that the experiential model yielded more brain-aligned word representations than taxonomic and distributional models; additionally, it contributed unique semantic information not represented by the other models.
@@ 3.1 Data and experiential model
For our experiments, we used word stimuli, fMRI responses and experiential model from Study 2 by Fernandino et al. (2022). [^1] These materials have been made publicly available by Fernandino et al. The full list of words and the experiential features can be found at https://www.pnas.org/doi/10.1073/pnas.2108091119#supplementary-materials; fMRI data are available at https://osf.io/87chb/. We describe each below.
@@
fMRI responses were collected from $36$ participants while they viewed the above-mentioned word stimuli one at a time and rated the frequency with which they experienced their corresponding entities in daily life. We focused on voxels from a ‘semantic network ROI’ defined by Binder et al. (2009) based on a meta-analysis and averaged responses across participants.
@@ 3.2 Computational models
Our model choices were motivated by the goal to maximise comparability across architectures. More concretely, we selected three models (language-only, vision-language, and audio-language) comparable in terms of fine-tuning objective—the contrastive one—and architecture—they all have a pretrained BERT Devlin et al. (2019) as language encoder. [^2] All three models were released with both BERT-based Devlin et al. (2019) and RoBERTa-based Liu et al. (2019) implementations. We used the former in all our experiments. One aspect in which these architectures differ is the amount of training data; however, we believe this issue does not invalidate our results and further discuss it in Limitations (Appendix A).
@@
(Multimodal Contrastive Learning of Sentence Embeddings, Zhang et al., 2022) is a vision-language sentence encoder fine-tuned by jointly optimising a SimCSE objective and a CLIP-like Radford et al. (2021) objective. The fine-tuning data for the first objective was the same as SimCSE’s; as for the CLIP-like objective, where a matching pair was defined by an image and its caption, the fine-tuning data consisted of 83K images from MS-COCO Lin et al. (2014) annotated with multiple captions.
@@
(Contrastive Language Audio Pretraining, Wu et al., 2023) is an audio-language model whose language encoder was initialised with pre-trained BERT weights and fine-tuned on audio-caption pairs with a CLIP-like objective. The fine-tuning data included $633,526$ audio-text pairs, with audio clips representing human activities, natural sounds, and audio effects.
@@
Devlin et al. (2019) is a transformer-based language-only model pretrained with two objectives: masked language modelling and next sentence prediction. Its pretraining data included the BooksCorpus (800M words, Zhu et al., 2015) and English Wikipedia (2500 words). As mentioned above, SimCSE, MCSE and CLAP fine-tuned pretrained BERT architectures.
@@
Li et al. (2019) is a vision-language model consisting of a BERT-based language encoder (initialised with parameters from pretrained BERT) and a pretrained visual feature extractor based on Faster RCNN (Ren et al., 2015). Its training objectives, which mirror BERT’s, were masked language modelling with image input and sentence-image prediction. The vision-language pretraining data comprised MS-COCO and VQA 2.0 Goyal et al. (2017). Note that this is not a contrastive model; we included it for reference as it can be considered as a vision-language extension of BERT, but it is not directly comparable with MCSE, SimCSE and CLAP.
@@ 3.3 Extracting representations
Given that all the models we considered were trained to learn contextualised representations from sentences, single words may be an out-of-distribution input. Therefore, following an approach similar to May et al. (2019), we embedded the noun stimuli in a set of generic template sentences (e.g., Someone mentioned the [word], see Appendix B for the complete list) when passing them to the models. [^3] We empirically verified that passing words within templates rather than in isolation yields higher alignment with both the experiential model and brain responses (see Appendix C.1). For all templates, we derived word representations from the hidden states of each layer; more specifically, we selected the hidden states corresponding to the tokens of the target word and averaged them across templates.
@@ 3.4 Alignment evaluation
To compare model representations against Exp48 and brain responses, we used representational similarity analysis (RSA, Kriegeskorte et al., 2008), which quantifies the alignment between two representational spaces (either by two models or by a model and brain responses) as the correlation between representational dissimilarity matrices (RDMs). In our experiments, RDMs were populated with pairwise cosine distances between model representations or fMRI responses for all the unique word-pairs. The alignment between RDMs was calculated as a Spearman correlation.
RSA allows comparing models’ alignment with Exp48 or brain responses, but it does not reveal whether models explain shared variance or provide independent contributions. Fernandino et al. (2022) computed partial correlations to check how much brain-relevant information Exp48 shared with the other models they considered, i.e., two distributional models (Word2vec and GloVe; Mikolov et al., 2013; Pennington et al., 2014) and two taxonomic models (a WordNet-based model and a categorical one). We used the same approach to determine how much brain-relevant information our tested models share with Exp48 and with each other. Partial correlations were obtained by regressing the RDM from one model out of the RDM from another model and then computing a Spearman correlation between the residuals and the RDM from brain responses.
@@ 4.1 Exp48 and brain alignment across models
We performed RSA to obtain a first measure of model representations’ alignment with Exp48 and fMRI responses. This analysis was conducted on model representations averaged across the three layers yielding the highest alignment individually; note that these layers may differ when considering alignment to brain responses vs. Exp48 (see Appendix C.2 for a visualisation of layer-wise alignment). The results from RSA against brain responses and Exp48 are displayed in Figure 1. All Spearman correlations are statistically significant ($p$ < 0.05), except for CLAP’s correlation with brain responses ($p$ = 0.70); we additionally verified that all the pairwise differences between correlations are statistically significant. [^4] Statistical significance was determined by applying a Fisher transformation to the correlation coefficients from each pair of models and calculating the $p$-value associated with the difference between the two $z$-scores. All $p$-values were Bonferroni-corrected with $\alpha$ = 0.05. The same approach for verifying statistical significance was applied to all correlation comparisons throughout the paper.
@@ 5.1 Do caption-like templates result in improved brain alignment?
To control for this potential confound, we re-extracted word representations employing different templates whose structure was modelled around captions (e.g., There is an [object] in a [place], or A [person] is [verb in -ing] in a [place]). These structures were identified based on a manual inspection of captions from MS-COCO, which was part of both MCSE’s and VisualBERT’s training. Given the challenges of creating caption-like templates providing a fitting context for all the word stimuli, we used different sets of templates for each sub-category of words described in Section 3.1 (e.g., There is a [food-word] on a table in a restaurant or A few people gathered for a [social event-word]). We provide the complete list of templates in Appendix B.
@@
6 Discussion and Conclusions
Our study provides an in-depth comparison between multimodal and language-only architectures in terms of their ability to capture experiential semantic information and their alignment with brain responses. While multimodal models are often expected to learn additional semantic aspects that language-only models cannot learn, our results reveal that their word representations are less aligned with Exp48 and fMRI responses than those by LMs. Moreover, within multimodal models, the vision-language ones show moderate positive correlations with Exp48 and fMRI responses, while the audio-language one correlates weakly with Exp48 and does not yield a significant correlation with brain responses. A potential explanation for these counterintuitive findings is that the dimensions used to create Exp48 are moderately abstract, whereas the extra information learnt by multimodal models may concern lower-level features or patterns of co-occurrence. Alternatively, it could be that multimodal training is not as effective during fine-tuning as it would be in the pre-training stage. However, we verified that not even the language encoder from the powerful CLIP Radford et al. (2021)—pretrained contrastively on 400M image-text pairs and excluded from the main experiment as it is not directly comparable with the other architectures—yields more brain-aligned word representations than BERT and SimCSE (we report these results in Appendix C.3).

### v2

@@ Abstract
A common assumption in Computational Linguistics is that text representations learnt by multimodal models are richer and more human-like than those by language-only models, as they are grounded in images or audio—similar to how human language is grounded in real-world experiences. However, empirical studies checking whether this is true are largely lacking. We address this gap by comparing word representations from contrastive multimodal models vs. language-only ones in the extent to which they capture experiential information—as defined by an existing norm-based ‘experiential model’—and align with human fMRI responses. Our results indicate that, surprisingly, language-only models are superior to multimodal ones in both respects. Additionally, they learn more unique brain-relevant semantic information beyond that shared with the experiential model. Overall, our study highlights the need to develop computational models that better integrate the complementary semantic information provided by multimodal data sources.
https:/github.com/dmg-illc/exp-info-models-brain
@@ 1 Introduction
How to link language representations to the real-world entities they refer to is a long-standing issue within semantics—the ‘symbol-grounding problem’ Harnad (1990); Bender and Koller (2020). With the advent of large language models (LLMs) learning astounding linguistic abilities purely from text, this question has been reframed as the ‘vector-grounding problem’ Mollo and Millière (2023), gaining new relevance. While some researchers think that word meanings should be intended as purely symbolic Fodor (1983), others believe that words have meanings precisely because they are linked to specific entities, experiences or notions Barsalou (2008). Supporters of the latter view stress that human language acquisition is situated in a rich multimodal environment, where new words are learnt through interactions with objects and people Vigliocco et al. (2014). Theories of embodied cognition further highlight the importance of linking words to concrete experience not only for their acquisition but also for their comprehension. Indeed, according to these theories, understanding sentences involves engaging perceptual, motor or emotional simulations of their content (for an overview, see Kaschak et al., 2024).
@@
Operationalising the ‘extra-linguistic’ information that multimodal models are allegedly learning is a prerequisite for approaching this issue. We did this by relying on a semantic model introduced by Fernandino et al. (2022) to capture ‘experiential information’. This cognitive model represents words as $n$-dimensional arrays where each entry corresponds to aggregated human ratings on a pre-defined experiential attribute (e.g., Vision, Motion, Harm). We focused on a set of nouns and evaluated the alignment between their representations provided by the experiential model and those by comparable unimodal (language-only) and multimodal (vision-language and audio-language) computational models. This analysis allowed us to uncover if multimodal models indeed reflect more semantic information than language-only models. Next, we checked whether capturing experiential information translates into higher alignment with brain responses recorded with functional magnetic resonance imaging (fMRI) to the same set of nouns.
Our findings indicate several interesting trends. First, both vision-language and language-only models exhibit significant alignment with the experiential model and brain responses, while the audio-language model displays weak or non-significant correlations. Second, this alignment is more pronounced for language-only models, which appear to capture a great deal of brain-relevant information beyond experiential. Lastly, language-only models remain more brain-aligned than vision-language models even when focusing on a set of more concrete words, although the gap is reduced. Overall, our study shows that current multimodal models learn less brain-relevant information—both experiential and beyond—than comparable language-only models, highlighting the need to explore different approaches to construct multimodal word representations.
Figure 1: Overview of our experimental setup. Representations for the word stimuli are derived from three different sources: 1) fMRI responses; 2) multimodal and language-only contrastive models; 3) human ratings along the experiential dimensions of the Exp48 model. Next, pairwise distances between these word representations are used to populate representational dissimilarity matrices (RDMs). Finally, alignment between representational spaces is computed by correlating the off-diagonal elements of the RDMs.
@@ 2.1 Embodied cognition
For example, a series of works on the Action-sentence Compatibility Effect (ACE, Glenberg and Kaschak, 2002) and its subsequent variants Borreggine and Kaschak (2006); Zwaan and Taylor (2006); Bub and Masson (2012) revealed a significant difference in reaction times—attributed to motor simulations—when participants had to respond to a sentence (e.g., You passed the note to Art) with a movement matching (extending their arm) vs.  non-matching (retreating their arm) that mentioned in the sentence. Similarly, the sentence-picture verification task Stanfield and Zwaan (2001), where participants have to respond to a picture that is either compatible (an eagle with its wings outstretched) or incompatible (an eagle with its wings folded) with a sentence (The eagle is in the sky), and its variations Connell (2007); Hoeben Mannaert et al. (2017) have also been widely used to demonstrate the occurrence of perceptual simulation during language comprehension. In parallel, a line of neuroscientific studies have found evidence that semantic processing may activate motor (among others, Hauk et al., 2004; Tettamanti et al., 2005; Aziz-Zadeh et al., 2006) and perceptual brain regions Kiefer et al. (2008); Van Dam et al. (2012).
@@ 2.2 Multimodal models of semantics
Embodied cognition and related ideas, such as visual grounding, have percolated from Cognitive Science to Computational Linguistics, motivating attempts to build semantic models that learn representations from data sources beyond text. Early efforts in this direction (e.g., Bruni et al., 2014; Kiela and Bottou, 2014; Lazaridou et al., 2015; Silberer and Lapata, 2012, 2014) were characterised by 1) a focus on developing human-aligned computational models of meaning and 2) limited computational modelling resources (large datasets of paired image-text inputs did not exist at the time, nor did large transformer-based architectures).
Recently, multimodal models have become more powerful and found application on a variety of downstream tasks (e.g., image captioning, image retrieval, or visual question answering). Some seminal works used a contrastive objective to learn aligned image and text representations Radford et al. (2021); Jia et al. (2021), while others—often inspired by BERT’s Devlin et al. (2019) successes in language modelling—applied its underlying intuitions to the vision-language domain Tan and Bansal (2019); Li et al. (2019); Lu et al. (2019); Chen et al. (2020). Finally, state-of-the-art large vision-language models (LVLMs, e.g., Li et al., 2023; Wang et al., 2024; Liu et al., 2024; Deitke et al., 2024; Laurençon et al., 2024), usually combining a large language model (LLM) with an image encoder, can engage in strikingly human-like conversations about images. In contrast to the early attempts at multimodal modelling, these works share 1) a focus on solving, or improving performance on, downstream tasks, and 2) the availability of massive datasets and large models with billions of parameters.
@@ 2.3 Experiential models of semantics
Recently, a few approaches motivated by embodied cognition have introduced models of semantics aimed at capturing ‘experiential information’, i.e., aspects of meaning related to how humans ground language in experiences. These experiential models were constructed by asking human annotators to rank words on a set of pre-defined dimensions. For example, Fernandino et al. (2022) introduced an experiential model based on 48 dimensions spanning perceptual, emotional, and action-related constructs. In two fMRI studies, they found that the experiential model yields more brain-aligned word representations than taxonomic and distributional models; additionally, it contributes unique semantic information not represented by the other models.
@@ 3 Methods
A schematic of our experimental pipeline is provided in Figure 1. In the following, we describe in detail the word stimuli, brain responses, computational models and evaluation procedures.
@@ 3.1 Data and experiential model
For our experiments, we used word stimuli, fMRI responses and experiential model from Study 2 by Fernandino et al. (2022). [^1] These materials have been made publicly available by Fernandino et al. The full list of words and the experiential features can be found at https:/www.pnas.org/doi/10.1073/pnas.2108091119#supplementary-materials; fMRI data are available at https:/osf.io/87chb/. We describe each below.
@@
fMRI responses were collected from $36$ participants. While viewing the above-mentioned word stimuli one at a time, they were instructed to rate the frequency with which they experienced their corresponding entities in daily life. Voxel-wise activations (beta maps) for each noun relative to the mean signal across other nouns were estimated using linear regressions (for additional details, see Fernandino et al., 2022). Here, we focus on the betas from voxels within a ‘semantic network ROI’ defined by Binder et al. (2009) based on a meta-analysis. Voxel-wise beta coefficients can be arranged in vectors representing the brain response elicited by each noun.
@@ 3.2 Computational models
Our model choices were motivated by the goal to maximise comparability across architectures. More concretely, we selected three models (language-only, vision-language, and audio-language) comparable in terms of fine-tuning objective—the contrastive one—and architecture—they all have a pretrained BERT Devlin et al. (2019) as language encoder. [^2] All three models were released with both BERT-based Devlin et al. (2019) and RoBERTa-based Liu et al. (2019) implementations. We used the former in all our experiments. One aspect in which these architectures differ is the amount of training data; however, we believe this issue does not invalidate our results and further discuss it in Section 6.
@@
(Multimodal Contrastive Learning of Sentence Embeddings, Zhang et al., 2022) is a vision-language sentence encoder fine-tuned by jointly optimising a SimCSE objective and a CLIP-like Radford et al. (2021) objective. The fine-tuning data for the first objective is the same as SimCSE’s; as for the CLIP-like objective, where a matching pair was defined by an image and its caption, the fine-tuning data consists of 83K images from MS-COCO Lin et al. (2014) annotated with multiple captions.
@@
(Contrastive Language Audio Pretraining, Wu et al., 2023) is an audio-language model whose language encoder was initialised with pre-trained BERT weights and fine-tuned on audio-caption pairs with a CLIP-like objective. The fine-tuning data includes $633,526$ audio-text pairs, with audio clips representing human activities, natural sounds, and audio effects.
@@
Devlin et al. (2019) is a transformer-based language-only model pretrained with two objectives: masked language modelling and next sentence prediction. Its pretraining data includes the BooksCorpus (800M words, Zhu et al., 2015) and English Wikipedia (2500 words). As mentioned above, SimCSE, MCSE and CLAP fine-tuned pretrained BERT architectures.
@@
Li et al. (2019) is a vision-language model consisting of a BERT-based language encoder (initialised with parameters from pretrained BERT) and a pretrained visual feature extractor based on Faster RCNN (Ren et al., 2015). Its training objectives, which mirror BERT’s, were masked language modelling with image input and sentence-image prediction. The vision-language pretraining data comprises MS-COCO and VQA 2.0 Goyal et al. (2017). Note that this is not a contrastive model; we included it for reference as it can be considered as a vision-language extension of BERT, but it is not directly comparable with MCSE, SimCSE and CLAP.
@@ 3.3 Extracting representations
Given that all the models we considered were trained to learn contextualised representations from sentences, single words may be an out-of-distribution input. Therefore, following an approach similar to May et al. (2019), we embedded the noun stimuli in a set of generic template sentences (e.g., Someone mentioned the <word>, see Appendix A for the complete list) when passing them to the models. [^3] We empirically verified that passing words within templates rather than in isolation yields higher alignment with both the experiential model and brain responses (see Appendix B.1). For all templates, we derived word representations from the hidden states of each layer; more specifically, we selected the hidden states corresponding to the tokens of the target word and averaged them across templates.
@@ 3.4 Alignment evaluation
RSA
To compare model representations against Exp48 and brain responses, we used representational similarity analysis (RSA, Kriegeskorte et al., 2008), which quantifies the alignment between two representational spaces (either by two models or by a model and brain responses) as the correlation between representational dissimilarity matrices (RDMs). In our experiments, RDMs were populated with pairwise cosine distances between model representations or fMRI responses for all the unique word-pairs. fMRI RDMs for individual participants were averaged into one aggregated RDM. The alignment between this fMRI RDM and morel-derived RDMs was calculated as a Spearman correlation.
Partial correlations
While RSA allows comparing models’ alignment with Exp48 or brain responses, it does not reveal whether models explain shared variance or provide independent contributions. Fernandino et al. (2022) computed partial correlations to check how much brain-relevant information Exp48 shared with the other models they considered, i.e., two distributional models (Word2vec and GloVe; Mikolov et al., 2013; Pennington et al., 2014) and two taxonomic models (a WordNet-based model and a categorical one). We used the same approach to determine how much brain-relevant information our tested models share with Exp48 and with each other. Formally, partial correlations can be defined as follows: Consider the RDM from Model A $y$, the RDM from Model B $x$, and the RDM of the brain responses $z$. The partial correlation of Model A without Model B is $\rho(r_{i},z_{i})$, where $r_{i}=y_{i}-\hat{y}_{i}$ are the residuals from the linear regression with equation $\hat{y}_{i}=a+bx_{i}$.
@@ 4.1 Exp48 and brain alignment across models
We performed RSA to obtain a first measure of model representations’ alignment with Exp48 and fMRI responses. This analysis was conducted on model representations averaged across the three layers yielding the highest alignment individually; note that these layers may differ when considering alignment to brain responses vs. Exp48 (see Appendix B.2 for a visualisation of layer-wise alignment). The results from RSA against brain responses and Exp48 are displayed in Figure 2. All Spearman correlations are statistically significant ($p$ < 0.05), except for CLAP’s correlation with brain responses ($p$ = 0.70); we additionally verified that all the pairwise differences between correlations are statistically significant. [^4] Statistical significance was determined by applying a Fisher transformation to the correlation coefficients from each pair of models and calculating the $p$-value associated with the difference between the two $z$-scores. All $p$-values were Bonferroni-corrected with $\alpha$ = 0.05. The same approach for verifying statistical significance was applied to all correlation comparisons throughout the paper.
@@ 5.1 Do caption-like templates result in improved brain alignment?
To control for this potential confound, we re-extracted word representations employing different templates, whose structure was modelled around captions (e.g., There is an <object> in a <place>, or A <person> is <verb in -ing> in a <place>). These structures were identified based on a manual inspection of captions from MS-COCO, which was part of both MCSE’s and VisualBERT’s training. Given the challenges of creating caption-like templates providing a fitting context for all the word stimuli, we used different sets of templates for each sub-category of words described in Section 3.1 (e.g., There is a <food-word> on a table in a restaurant or A few people gathered for a <social event-word>). We provide the complete list of templates in Appendix A.
@@
6 Discussion
While multimodal models are often expected to incorporate additional semantic aspects that language-only models cannot learn, our results reveal that their word representations are less aligned with Exp48 and fMRI responses than those by LMs. Moreover, within multimodal models, the vision-language ones show moderate positive correlations with Exp48 and fMRI responses, while the audio-language one correlates weakly with Exp48 and does not yield a significant correlation with brain responses. Below, we discuss factors that may have played a role in these partially unexpected results.
Amount of training data
While being comparable in terms of learning objectives and architecture, SimCSE, MCSE and CLAP still differ in the amount of fine-tuning data. For the SimCSE–MCSE comparison, this does not appear to be a confound: Despite being fine-tuned on less sentences than MCSE, SimCSE still proves to be more Exp48- and brain-aligned. A potential reason for this may be that the grounded language employed in image captions causes a shift of semantic representations towards more concrete meanings. As for CLAP, the smaller amount of fine-tuning audio-caption pairs, together with its optimising only a CLIP-like objective (without a SimCSE-like one), may have played a role in its poor alignment.
Multimodal pretraining vs. fine-tuning
A potential explanation for the inferior performance of multimodal models could be that training on multimodal pairs is not as effective during fine-tuning as it is during pre-training. However, we verified that not even the language encoder from the powerful CLIP Radford et al. (2021) [^7] This model was excluded from the main experiment as it is not directly comparable with the other architectures.—pretrained contrastively on 400M image-text pairs—yields more brain-aligned word representations than BERT and SimCSE (see Appendix B.3).
Models from present vs. past studies